# Multi-Objective Agentic Rewrites for Unstructured Data Processing

### v1

@@ 1  Introduction
LLMs are now integrated into systems that support queries over unstructured data, from both industry and academia (databricks-llm; snowflake-llm; HeSethi2025BigQueryAIFunctions; shankar2025docetl; patel2025semantic; liu2024declarative; jo2024thalamusdb). In these LLM-powered data processing systems, a query is expressed as a sequence of semantic operators. Semantic operators are data processing operators such as map, reduce, and filter, each described in natural language for an LLM to carry out. Users define an initial pipeline, and the system’s optimizer then determines how to execute it. DocETL (shankar2025docetl) is one such system. Consider the following example of a DocETL workload from the public defender’s office in a major California city:
@@
LOTUS (patel2025semantic) No search
ABACUS (russo2025abacus)
DocETL-V1 (shankar2025docetl)
@@
Limitations of Existing Systems. Query optimizers in data systems typically rely on the principle of optimal substructure: an optimal plan can be constructed from optimal solutions to its subplans. [^1] Even when physical properties like sort order make subplan costs context-dependent, we can restore optimal substructure, by treating each (expression, property) pair as a distinct subproblem. For example, the Cascades framework (Graefe1995TheCF) organizes subexpressions into equivalence groups, and the optimal plan for a subexpression is reused wherever it appears. However, in LLM-powered data processing, even when optimizing only for accuracy, the same subplan can produce outputs of varying benefit, depending on the subplans that precede or follow it. For instance, in Ex. 1.1, suppose we decompose the pipeline into a map per factor type, followed by a reduce that unifies the evidence. A map that extracts smaller or fewer text spans may achieve higher accuracy in isolation, but a map that includes more surrounding context may yield better overall accuracy if the downstream reduce can leverage that context to deduplicate extractions or filter out false positives. Which map implementation is optimal depends on which reduce implementation that follows—so we cannot optimize them independently. More generally, since individual subplan choices interact in unpredictable ways, composing locally optimal subplans can yield globally suboptimal plans.
Existing optimizers for LLM-powered data processing have limitations as summarized in Table 1. LOTUS (patel2025semantic) only minimizes cost, requiring users to author an accurate plan first, which is difficult for users. LOTUS ignores the pipeline search problem entirely by considering only one optimized implementation per semantic operator. ABACUS (russo2025abacus) performs cost-based search using transformation rules (i.e., logical-to-logical) from traditional data processing (e.g., filter pushdown) and implementation rules (i.e., logical-to-physical) for semantic operators (e.g., model selection, prompting strategies, ensembling). ABACUS intelligently samples operator implementations to estimate their cost and accuracy, then uses the Cascades framework (Graefe1995TheCF), which relies on optimal substructure to search for Pareto-optimal query plans. Our original DocETL optimizer (shankar2025docetl) (that we call DocETL-V1) relaxes the optimal substructure assumption slightly in its search algorithm, but still optimizes operators from upstream to downstream, potentially missing beneficial rewrites where changing downstream operators may benefit from different upstream choices. Moreover, DocETL-V1 optimizes only for accuracy, not cost.
@@
Then, during the rewrite process, we observed that when the LLM agent receives the full specifications of all 30+ directives simultaneously in its prompt, it struggles to select appropriate rewrites. Inspired by progressive disclosure from HCI (CarrollRosson1987ProgressiveDisclosure)—a technique that reduces cognitive load in user interfaces by revealing information gradually—we structure the agent’s interaction with directives in stages. Agents initially see only directive names and high-level descriptions. When an agent selects a directive, it loads the full specification (detailed descriptions, instantiation schemas, examples) on-demand. Additionally, the agent can invoke a read_document tool at any point during instantiation to inspect sample data, grounding its decisions in actual document characteristics.
@@
New Search Algorithm. Another challenge is designing an efficient global search algorithm. Since evaluating a pipeline requires executing it on sample data, we can only explore a limited number of pipelines. Rewrites can be applied in sequence, creating a vast search space, as shown in Fig. 1. For instance, the highest-accuracy pipeline (#46, green) is discovered through four rewrites: first switching to a cheaper model (#2), then applying data decomposition to process document chunks (#6), followed by prompt improvements to the map (#24) and reduce (#46) operations. While LLM agents can determine which directives to apply and how to instantiate them, it is unclear which pipelines will lead to high-accuracy, low-cost descendants. Our insight is to learn which pipelines are promising to rewrite, by adapting multi-armed bandits. We use a variant of the UCB (Upper Confidence Bound) algorithm, adapted for the tree setting (auer2002finite), to select which pipeline to rewrite. However, we define a custom metric rather than using the typical hypervolume metric from multi-objective optimization (zitzler1999evolutionary), which treats all Pareto frontier points as equally valuable. In LLM-powered data processing, low-accuracy pipelines vastly outnumber high-accuracy ones, so optimizing for hypervolume would waste the evaluation budget exploring low-accuracy regions of the search space. We instead introduce a metric based on marginal accuracy contribution—the vertical distance between a pipeline’s accuracy and the best accuracy at comparable cost (shown as the red line in Fig. 3)—and score each pipeline by aggregating this metric across the pipeline and its descendants.
@@
Then, with over 30 directives, each applicable to multiple operators, a single pipeline could spawn hundreds of children, exhausting the evaluation budget on variants of one pipeline rather than exploring deeper rewrite sequences. We employ progressive widening (chaslot2008progressive), a technique that limits how many edges (i.e., immediate rewrites) a node (i.e., pipeline) can have based on its visit count (i.e., total number of descendants of any depth). As a node accumulates more visits, it is allowed to have more edges, but its edge growth is sublinear, forcing the search to explore other regions of the graph before returning to generate additional variants from any single pipeline.
@@ 2  Background and Definitions
We build our optimizer on top of the DocETL (shankar2025docetl) system (i.e., DSL, parser, and execution engine), though our techniques can extend to other systems. We write DocETL-V1 when referring to contributions from shankar2025docetl, such as the original query optimizer and rewrite directives. Secs. 2.1 and 2.2 describe background on DocETL and semantic operators. Sec. 2.3 describes the optimization problem setup. Table 3 summarizes all notation.
@@ 2.1. Datasets, Operators, and Pipelines
Semantic Operators. We refer to operators that transform data using natural language (NL) specifications as semantic operators, following patel2025semantic. Each semantic operator in DocETL has four components: (i) an operator type such as map, filter, or reduce (e.g., map applies a transformation to each document, reduce aggregates groups of documents into one); (ii) a prompt template written in natural language that describes the operator’s semantics, expressed in Jinja (jinja2); (iii) an output schema that declares the structure of the operator’s results; and (iv) a model specifying the LLM used to execute the operator (e.g., gpt-4o-mini). We denote $M$ as the set of models that the user makes available to the optimizer, such as GPT or Gemini variants, and by $m\in M$ a particular model. While the user selects a specific $m$ when authoring an operator, the optimizer is free to choose any model from $M$ for any semantic operator, or replace LLM execution with code-powered implementations—synthesized code that realizes the task specified by the operator’s prompt template and output schema.
@@
Pipelines. Given operators, either LLM or code-powered, a pipeline $P$ is then a sequence of $k$ semantic operators $(o^{(1)},\ldots,o^{(k)})$; given an input dataset $D$, we write its execution as $o^{(1)}\to o^{(2)}\to\cdots\to o^{(k)}$, where $\to$ denotes function composition, following shankar2025docetl: $o^{(k)}\big(o^{(k-1)}(\cdots o^{(1)}(D)\cdots)\big)$. While users specify operators via NL prompt templates, schemas, and model choices, these serve as a baseline specification. The optimizer may rewrite pipelines by substituting models or replacing LLM-powered implementations with code-powered ones (including relational operators), decomposing operators into multiple ones, or fusing multiple operators, to discover cheaper or more accurate alternatives. DocETL-V1 supports six semantic operator types (map, parallel_map, reduce, filter, resolve, equijoin) and three operators (split, gather, unnest). The auxiliary operators do not invoke LLMs. MOAR extends this with additional operator types detailed in Sec. 3. In DocETL, pipelines and operators are specified in YAML (yaml12): a pipeline is a list of operator configurations, where each operator’s configuration is represented as a dictionary of parameters (such as prompt template, output schema, and model).
@@ 2.3. Optimization Problem Setup
17 Few-shot Examples ${\scriptsize\texttt{o}}_{x}\Rightarrow{\color[rgb]{0.046875,0.5078125,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.046875,0.5078125,0.3515625}{\scriptsize\texttt{o}}_{x^{\prime}}}$ where $x^{\prime}=({\color[rgb]{0.046875,0.5078125,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.046875,0.5078125,0.3515625}p^{\prime}},s,m)$ Adds few-shot examples to prompts (a standard strategy for improving accuracy (brown2020language)), thus making the task “easier” for the LLM. [Uncaptioned image]
@@ 3  New Operators and Directives
DocETL-V1 proposed three categories of rewrite directives targeting accuracy: projection synthesis (i.e., creating map operators that decompose tasks), data decomposition (i.e., splitting documents into chunks or groups of many documents into smaller-size groups), and LLM-centric improvements (e.g., prompting strategies) (shankar2025docetl). However, these directives typically increase cost. MOAR extends DocETL’s directive library in two ways. First, we introduce two new categories: fusion and reordering (combining or rearranging operators) and code synthesis (replacing or adding code-powered operations). Second, we adapt existing categories to also reduce cost—e.g., projection synthesis can compress documents to reduce tokens processed by downstream operators, data decomposition can limit processing to relevant chunks, and LLM-centric rewrites can substitute cheaper models. Together with DocETL-V1’s 13 directives, MOAR’s 18 new directives bring the total to over 30. Table 2 summarizes the new directives; detailed descriptions are in App. B. We highlight a few examples below (directive numbers reference the corresponding row in Table 2):
@@
To support the new directives, we introduce three new operator types. The sample operator selects a subset of documents (or chunks) most relevant to the downstream operator. It may use BM25 keyword search (robertson2009probabilistic), embedding-based similarity, random sampling, or stratified variants of these methods that ensure each subgroup (e.g., based on metadata keys in the document) is proportionally represented in the sample. For instance, when extracting enhancement factors from police reports (Ex. 1.1), if a prior split operator has divided each report into chunks—each chunk now becoming a document—then sample can issue a BM25 query with terms like “firearm” and “weapon” or use embedding similarity to a query such as “threatening with a weapon” to select only the relevant chunks before the downstream map. By processing fewer documents, the subsequent LLM-powered operator incurs lower cost. Next, the extract operator presents a document’s JSON representation with line numbers to an LLM, which returns ranges relevant to the operator’s natural language specification (e.g., “lines 45–67”); only those lines are retained, preserving the document’s key–value structure. Unlike a map that outputs text verbatim, extract guarantees exact subsets of documents and requires far fewer output tokens (thus reducing cost). Finally, code-powered operators (code_map, code_reduce, code_filter) execute synthesized Python instead of invoking an LLM. Overall, the aforementioned operators enable directives to reduce costs by processing smaller document portions (sample, extract) or replacing LLM calls with code. A complete operator list appears in Table 7.
@@ 4  Search Algorithm
Search Space Representation. Unlike traditional query optimizers that construct plans from optimal subplans (selinger1979access; Graefe1995TheCF), MOAR performs global search over complete pipelines. This design reflects our observation from Sec. 1 that the best plan for a semantic operator pipeline may rely on subplans that are individually suboptimal, because the benefit of a subplan isn’t independent of its outputs. To support this global search, we represent the search space as a tree $T=(V,E)$ with the user-authored pipeline $P_{0}$ at the root. Each node $P\in V$ represents a complete pipeline configuration. Each edge $P\xrightarrow{r}P^{\prime}$ applies a single rewrite $r$ to produce child pipeline $P^{\prime}$. Every node has exactly one parent and may have multiple children (shown as the search tree in Fig. 3). The path from $P_{0}$ to any node captures the sequence of rewrites used to construct it.
@@ 7     $P^{\star}\leftarrow$ Select($P_{0},T_{t}$);
Algorithm Overview. MOAR searches for better pipelines by iteratively selecting, rewriting, and evaluating candidates, as described in Algorithm 1 and illustrated in Fig. 3. The search process begins with initialization, where MOAR constructs an initial frontier that provides diverse starting points across the accuracy–cost spectrum. After initialization, each iteration of the search loop proceeds through two phases. First, in the selection phase, MOAR adapts a multi-armed bandit framework (auer2002finite) to decide which pipeline to rewrite, guided by how much the pipeline and its descendants contribute to the Pareto frontier. Second, in the rewriting and evaluation phase, an LLM agent chooses and instantiates a rewrite directive to produce a child pipeline. The child $P^{\prime}$ is then executed on $D_{o}$ to obtain $(\hat{c}(P^{\prime}),\hat{a}(P^{\prime}))$, which informs selection in future iterations. Each iteration incurs significant latency due to LLM-guided rewriting and pipeline execution. MOAR therefore parallelizes search across multiple workers, with only the selection phase synchronized to ensure consistency. The loop repeats until the evaluation budget $B$ is exhausted, returning the final frontier $\hat{\mathcal{F}}_{B}$.

### v3

@@ 1  Introduction
LLMs are now integrated into systems that support queries over unstructured data, from both industry and academia (Wendell et al., 2023; Snowflake, 2025; He and Sethi, 2025; Shankar et al., 2025a; Patel et al., 2025; Liu et al., 2025a; Jo and Trummer, 2024). In these LLM-powered data processing systems, a query is expressed as a sequence of semantic operators. Semantic operators are data processing operators such as map, reduce, and filter, each described in natural language for an LLM to carry out. Users define an initial pipeline, and the system’s optimizer then determines how to execute it. DocETL (Shankar et al., 2025a) is one such system. Consider the following example of a DocETL workload from the public defender’s office in a major California city:
@@
LOTUS (Patel et al., 2025) No search
ABACUS (Russo et al., 2025)
DocETL-V1 (Shankar et al., 2025a)
@@
Limitations of Existing Systems. Query optimizers in data systems typically rely on the principle of optimal substructure: an optimal plan can be constructed from optimal solutions to its subplans. [^1] Even when physical properties like sort order make subplan costs context-dependent, we can restore optimal substructure, by treating each (expression, property) pair as a distinct subproblem. For example, the Cascades framework (Graefe, 1995) organizes subexpressions into equivalence groups, and the optimal plan for a subexpression is reused wherever it appears. However, in LLM-powered data processing, even when optimizing only for accuracy, the same subplan can produce outputs of varying benefit, depending on the subplans that precede or follow it. For instance, in Ex. 1.1, suppose we decompose the pipeline into a map per factor type, followed by a reduce that unifies the evidence. A map that extracts smaller or fewer text spans may achieve higher accuracy in isolation, but a map that includes more surrounding context may yield better overall accuracy if the downstream reduce can leverage that context to deduplicate extractions or filter out false positives. Which map implementation is optimal depends on which reduce implementation that follows—so we cannot optimize them independently. More generally, since individual subplan choices interact in unpredictable ways, composing locally optimal subplans can yield globally suboptimal plans.
Existing optimizers for LLM-powered data processing have limitations as summarized in Table 1. LOTUS (Patel et al., 2025) only minimizes cost, requiring users to author an accurate plan first, which is difficult for users. LOTUS ignores the pipeline search problem entirely by considering only one optimized implementation per semantic operator. ABACUS (Russo et al., 2025) performs cost-based search using transformation rules (i.e., logical-to-logical) from traditional data processing (e.g., filter pushdown) and implementation rules (i.e., logical-to-physical) for semantic operators (e.g., model selection, prompting strategies, ensembling). ABACUS intelligently samples operator implementations to estimate their cost and accuracy, then uses the Cascades framework (Graefe, 1995), which relies on optimal substructure to search for Pareto-optimal query plans. Our original DocETL optimizer (Shankar et al., 2025a) (that we call DocETL-V1) relaxes the optimal substructure assumption slightly in its search algorithm, but still optimizes operators from upstream to downstream, potentially missing beneficial rewrites where changing downstream operators may benefit from different upstream choices. Moreover, DocETL-V1 optimizes only for accuracy, not cost.
@@
Then, during the rewrite process, we observed that when the LLM agent receives the full specifications of all 30+ directives simultaneously in its prompt, it struggles to select appropriate rewrites. Inspired by progressive disclosure from HCI (Carroll and Rosson, 1987)—a technique that reduces cognitive load in user interfaces by revealing information gradually—we structure the agent’s interaction with directives in stages. Agents initially see only directive names and high-level descriptions. When an agent selects a directive, it loads the full specification (detailed descriptions, instantiation schemas, examples) on-demand. Additionally, the agent can invoke a read_document tool at any point during instantiation to inspect sample data, grounding its decisions in actual document characteristics.
@@
New Search Algorithm. Another challenge is designing an efficient global search algorithm. Since evaluating a pipeline requires executing it on sample data, we can only explore a limited number of pipelines. Rewrites can be applied in sequence, creating a vast search space, as shown in Fig. 1. For instance, the highest-accuracy pipeline (#46, green) is discovered through four rewrites: first switching to a cheaper model (#2), then applying data decomposition to process document chunks (#6), followed by prompt improvements to the map (#24) and reduce (#46) operations. While LLM agents can determine which directives to apply and how to instantiate them, it is unclear which pipelines will lead to high-accuracy, low-cost descendants. Our insight is to learn which pipelines are promising to rewrite, by adapting multi-armed bandits. We use a variant of the UCB (Upper Confidence Bound) algorithm, adapted for the tree setting (Auer et al., 2002), to select which pipeline to rewrite. However, we define a custom metric rather than using the typical hypervolume metric from multi-objective optimization (Zitzler, 1999), which treats all Pareto frontier points as equally valuable. In LLM-powered data processing, low-accuracy pipelines vastly outnumber high-accuracy ones, so optimizing for hypervolume would waste the evaluation budget exploring low-accuracy regions of the search space. We instead introduce a metric based on marginal accuracy contribution—the vertical distance between a pipeline’s accuracy and the best accuracy at comparable cost (shown as the red line in Fig. 3)—and score each pipeline by aggregating this metric across the pipeline and its descendants.
@@
Then, with over 30 directives, each applicable to multiple operators, a single pipeline could spawn hundreds of children, exhausting the evaluation budget on variants of one pipeline rather than exploring deeper rewrite sequences. We employ progressive widening (Chaslot et al., 2008), a technique that limits how many edges (i.e., immediate rewrites) a node (i.e., pipeline) can have based on its visit count (i.e., total number of descendants of any depth). As a node accumulates more visits, it is allowed to have more edges, but its edge growth is sublinear, forcing the search to explore other regions of the graph before returning to generate additional variants from any single pipeline.
@@ 2  Background and Definitions
We build our optimizer on top of the DocETL (Shankar et al., 2025a) system (i.e., DSL, parser, and execution engine), though our techniques can extend to other systems. We write DocETL-V1 when referring to contributions from Shankar et al. (2025a), such as the original query optimizer and rewrite directives. Secs. 2.1 and 2.2 describe background on DocETL and semantic operators. Sec. 2.3 describes the optimization problem setup. Table 3 summarizes all notation.
@@ 2.1. Datasets, Operators, and Pipelines
Semantic Operators. We refer to operators that transform data using natural language (NL) specifications as semantic operators, following Patel et al. (2025). Each semantic operator in DocETL has four components: (i) an operator type such as map, filter, or reduce (e.g., map applies a transformation to each document, reduce aggregates groups of documents into one); (ii) a prompt template written in natural language that describes the operator’s semantics, expressed in Jinja (Ronacher, 2008); (iii) an output schema that declares the structure of the operator’s results; and (iv) a model specifying the LLM used to execute the operator (e.g., gpt-4o-mini). We denote $M$ as the set of models that the user makes available to the optimizer, such as GPT or Gemini variants, and by $m\in M$ a particular model. While the user selects a specific $m$ when authoring an operator, the optimizer is free to choose any model from $M$ for any semantic operator, or replace LLM execution with code-powered implementations—synthesized code that realizes the task specified by the operator’s prompt template and output schema.
@@
Pipelines. Given operators, either LLM or code-powered, a pipeline $P$ is then a sequence of $k$ semantic operators $(o^{(1)},\ldots,o^{(k)})$; given an input dataset $D$, we write its execution as $o^{(1)}\to o^{(2)}\to\cdots\to o^{(k)}$, where $\to$ denotes function composition, following Shankar et al. (2025a): $o^{(k)}\big(o^{(k-1)}(\cdots o^{(1)}(D)\cdots)\big)$. While users specify operators via NL prompt templates, schemas, and model choices, these serve as a baseline specification. The optimizer may rewrite pipelines by substituting models or replacing LLM-powered implementations with code-powered ones (including relational operators), decomposing operators into multiple ones, or fusing multiple operators, to discover cheaper or more accurate alternatives. DocETL-V1 supports six semantic operator types (map, parallel_map, reduce, filter, resolve, equijoin) and three operators (split, gather, unnest). The auxiliary operators do not invoke LLMs. MOAR extends this with additional operator types detailed in Sec. 3. In DocETL, pipelines and operators are specified in YAML (Ben-Kiki et al., 2009): a pipeline is a list of operator configurations, where each operator’s configuration is represented as a dictionary of parameters (such as prompt template, output schema, and model).
@@ 2.3. Optimization Problem Setup
17 Few-shot Examples ${\scriptsize\texttt{o}}_{x}\Rightarrow{\color[rgb]{0.046875,0.5078125,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.046875,0.5078125,0.3515625}{\scriptsize\texttt{o}}_{x^{\prime}}}$ where $x^{\prime}=({\color[rgb]{0.046875,0.5078125,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.046875,0.5078125,0.3515625}p^{\prime}},s,m)$ Adds few-shot examples to prompts (a standard strategy for improving accuracy (Brown et al., 2020)), thus making the task “easier” for the LLM. [Uncaptioned image]
@@ 3  New Operators and Directives
DocETL-V1 proposed three categories of rewrite directives targeting accuracy: projection synthesis (i.e., creating map operators that decompose tasks), data decomposition (i.e., splitting documents into chunks or groups of many documents into smaller-size groups), and LLM-centric improvements (e.g., prompting strategies) (Shankar et al., 2025a). However, these directives typically increase cost. MOAR extends DocETL’s directive library in two ways. First, we introduce two new categories: fusion and reordering (combining or rearranging operators) and code synthesis (replacing or adding code-powered operations). Second, we adapt existing categories to also reduce cost—e.g., projection synthesis can compress documents to reduce tokens processed by downstream operators, data decomposition can limit processing to relevant chunks, and LLM-centric rewrites can substitute cheaper models. Together with DocETL-V1’s 13 directives, MOAR’s 18 new directives bring the total to over 30. Table 2 summarizes the new directives; detailed descriptions are in App. B. We highlight a few examples below (directive numbers reference the corresponding row in Table 2):
@@
To support the new directives, we introduce three new operator types. The sample operator selects a subset of documents (or chunks) most relevant to the downstream operator. It may use BM25 keyword search (Robertson et al., 2009), embedding-based similarity, random sampling, or stratified variants of these methods that ensure each subgroup (e.g., based on metadata keys in the document) is proportionally represented in the sample. For instance, when extracting enhancement factors from police reports (Ex. 1.1), if a prior split operator has divided each report into chunks—each chunk now becoming a document—then sample can issue a BM25 query with terms like “firearm” and “weapon” or use embedding similarity to a query such as “threatening with a weapon” to select only the relevant chunks before the downstream map. By processing fewer documents, the subsequent LLM-powered operator incurs lower cost. Next, the extract operator presents a document’s JSON representation with line numbers to an LLM, which returns ranges relevant to the operator’s natural language specification (e.g., “lines 45–67”); only those lines are retained, preserving the document’s key–value structure. Unlike a map that outputs text verbatim, extract guarantees exact subsets of documents and requires far fewer output tokens (thus reducing cost). Finally, code-powered operators (code_map, code_reduce, code_filter) execute synthesized Python instead of invoking an LLM. Overall, the aforementioned operators enable directives to reduce costs by processing smaller document portions (sample, extract) or replacing LLM calls with code. A complete operator list appears in Table 7.
@@ 4  Search Algorithm
Search Space Representation. Unlike traditional query optimizers that construct plans from optimal subplans (Selinger et al., 1979; Graefe, 1995), MOAR performs global search over complete pipelines. This design reflects our observation from Sec. 1 that the best plan for a semantic operator pipeline may rely on subplans that are individually suboptimal, because the benefit of a subplan isn’t independent of its outputs. To support this global search, we represent the search space as a tree $T=(V,E)$ with the user-authored pipeline $P_{0}$ at the root. Each node $P\in V$ represents a complete pipeline configuration. Each edge $P\xrightarrow{r}P^{\prime}$ applies a single rewrite $r$ to produce child pipeline $P^{\prime}$. Every node has exactly one parent and may have multiple children (shown as the search tree in Fig. 3). The path from $P_{0}$ to any node captures the sequence of rewrites used to construct it.
@@ 7     $P^{\star}\leftarrow$ Select($P_{0},T_{t}$);
Algorithm Overview. MOAR searches for better pipelines by iteratively selecting, rewriting, and evaluating candidates, as described in Algorithm 1 and illustrated in Fig. 3. The search process begins with initialization, where MOAR constructs an initial frontier that provides diverse starting points across the accuracy–cost spectrum. After initialization, each iteration of the search loop proceeds through two phases. First, in the selection phase, MOAR adapts a multi-armed bandit framework (Auer et al., 2002) to decide which pipeline to rewrite, guided by how much the pipeline and its descendants contribute to the Pareto frontier. Second, in the rewriting and evaluation phase, an LLM agent chooses and instantiates a rewrite directive to produce a child pipeline. The child $P^{\prime}$ is then executed on $D_{o}$ to obtain $(\hat{c}(P^{\prime}),\hat{a}(P^{\prime}))$, which informs selection in future iterations. Each iteration incurs significant latency due to LLM-guided rewriting and pipeline execution. MOAR therefore parallelizes search across multiple workers, with only the selection phase synchronized to ensure consistency. The loop repeats until the evaluation budget $B$ is exhausted, returning the final frontier $\hat{\mathcal{F}}_{B}$.